Grade 12th passPhysical Chemistry

For the estimation of nitrogen, 1.4 g of an organic compound was digested by Kjeldahl method and the evolved ammonia was absorbed in 60 mL of M10 sulphuric acid. The unreacted acid required 20 mL of M10 sodium hydroxide for complete neutralization. The percentage of nitrogen in the compound is?

Prasoon
9 Years agoGrade 12th pass

1 Answer

sourabh
9 Years ago
%N=141000×100×0.1×1001.4%N=141000×100×0.1×1001.4=10%=10%this is your answer that you ask ,thank you